OpenStudy (anonymous):

Does anyone know how to integrate this function? (2x + 2) / ((x^2 +1)( x - 1)^3) I think we're suppose to use partial fractions or something Mahalos

OpenStudy (anonymous):

yes u have to use partial functions

OpenStudy (anonymous):

\[\frac{2 x+2}{(x-1)^3 \left(x^2+1\right)}=\frac{1}{x^2+1}-\frac{1}{(x-1)^2}+\frac{2}{(x-1)^3} \]